# Eccentricity distribution of extreme mass ratio inspirals

### v1

@@ IV Comparison with previous works
Whilst a precise one-to-one comparison between the two calculations is difficult, the observed differences may be due to their assumption that the system reaches a steady state. In fact, we find that for $M_{\bullet}<10^{6}\,\mathrm{M_{\odot}}$ two-body relaxation continuously modifies the stellar distribution, so that both the EMRI formation rate and the stellar profiles are not steady (see the upper panel of Fig. 3). On the other hand, for $M_{\bullet}\geq 10^{6}\,\mathrm{M_{\odot}}$ it takes more than $10$ Gyr for the onset of a profile determined entirely by relaxation, as this is the time required to lose memory of initial conditions inside $R_{\mathrm{inf}}$ [86].
@@
Finally, Babak et al. [27] reported a flat eccentricity distribution at plunge in the range $0<e_{\mathrm{pl}}<0.2$, with a small tail of outliers with higher $e_{\mathrm{pl}}$. This distribution was derived as a suitable proxy to a set of numerical integrations of EMRIs within the classical capture radius $a_{\mathrm{c}}$ to plunge. It should be noted, however, that although the $e_{\mathrm{pl}}$ information is given in their paper, it is not used at any point in their calculations, which focus on the computation of cosmological EMRI rates rather than on the determination of detailed EMRI orbital properties. Nevertheless, a flat $e_{\mathrm{pl}}$ distribution in the range $[0,0.2]$ has become a somewhat standard assumption when simulating EMRI populations. Our work demonstrates how inaccurate this assumption is (see Fig. 4), and highlights the much larger eccentricity range covered by EMRIs formed via two-body relaxation.
@@ V Conclusions
In this work we derived realistic EMRI eccentricity distributions around Schwarzschild MBHs assuming the two-body relaxation channel. We quote the distribution at different stages of the inspiral, up to the final plunge. We accomplished this by integrating further the simulations performed by Mancieri et al. [67] using the FEW package, appropriately weighting each run proportionally to the EMRI formation rate at its initial semi-major axis.
@@
Our results call for an extension of the capabilities of FEW in order not to miss a large part of the EMRI population. We stress that any waveform model that aims at modeling astrophysical EMRIs has to cover a much wider parameter space that currently available. Accurate modeling of EMRIs is fundamental for parameter estimation, tests of general relativity and environmental effects, and for avoiding spurious residual in LISA global fit procedures. Precisely detecting high eccentricity EMRIs is also essential if we wish to infer the underling formation rate of each channel, which in turn will reveal more about the nature of MBHs and their host galaxies.
@@ Acknowledgements.
We thank Niels Warburton for useful discussions.
D.M. acknowledges that this publication was produced while attending the PhD program in Space Science and Technology at the University of Trento, Cycle XXXIX, with the support of a scholarship financed by the Ministerial Decree no. 118 of 2nd March 2023, based on the NRRP - funded by the European Union - NextGenerationEU - Mission 4 “Education and Research”, Component 1 “Enhancement of the offer of educational services: from nurseries to universities” - Investment 4.1 “Extension of the number of research doctorates and innovative doctorates for public administration and cultural heritage” - CUP E66E23000110001. A.S. acknowledges the financial support provided under the European Union’s H2020 ERC Advanced Grant “PINGU” (Grant Agreement: 101142079). M.B. acknowledges support from the Italian Ministry for Universities and Research (MUR) program “Dipartimenti di Eccellenza 2023-2027”, within the framework of the activities of the Centro Bicocca di Cosmologia Quantitativa (BiCoQ).
@@ Appendix A Alternative weighting
In this appendix, we examine how the results in Fig. 4 change when using an alternative weighting procedure based on the classical criterion for EMRI formation. In the standard approach [91], BHs entering the loss cone are classified as EMRIs if $a<a_{\mathrm{c}}$, and as direct plunges if $a>a_{\mathrm{c}}$. By contrast, in Sect. II.2 we employed a transfer function $S$ to provide a smoother transition between the two regimes and to account for cliffhanger EMRIs.
@@
in the energy range explored in Mancieri et al. [67]. We then re-process the same simulations as in Sect. II.2 to compute the formation rate of EMRIs according to the classical differential rate
@@
Figure 8: EMRI eccentricity distribution at plunge, for different central MBH masses. Here we report the probability density function of our EMRI population analyzed with the procedure described in Sect. II.2 (blue solid histograms) and the one described in this appendix (green dotted histograms). We also show (orange dashed histogram) the distribution reported in Babak et al. [27], which has become a standard assumption in the literature.

### v2

@@ IV Comparison with previous works
While a precise one-to-one comparison between the two calculations is difficult, the observed differences may be due to their assumption that the system reaches a steady state. In fact, we find that for $M_{\bullet}<10^{6}\,\mathrm{M_{\odot}}$ two-body relaxation continuously modifies the stellar distribution, so that both the EMRI formation rate and the stellar profiles are not steady (see Fig. 3 and the upper panel of Fig. 4). On the other hand, for $M_{\bullet}\geq 10^{6}\,\mathrm{M_{\odot}}$ it takes more than $10$ Gyr for the onset of a profile determined entirely by relaxation, as this is the time required to lose memory of initial conditions inside $R_{\mathrm{inf}}$ [2022MNRAS.514.3270B].
Around the same time our work was released, 2025arXiv250900469S reported EMRI eccentricity distributions for several cases, including the two-body formation channel. They evolve the energy and angular momentum distributions of stars and stellar-mass BHs in nuclear star clusters using a Fokker–Planck formalism, identify the boundary where $t_{\mathrm{rlx}}=t_{\mathrm{GW}}$, and transport the flux crossing it to orbits with pericenter $r_{\mathrm{p}}=10\,R_{\mathrm{g}}$. The resulting eccentricity distributions are consistently weighted through this flux, converting from energy and angular momentum to orbital parameters. Compared to our approach, theirs is inherently Newtonian and uses the GW emission equations from 1964PhRv..136.1224P. They also cannot account for cliffhanger EMRIs, as their method relies on orbit-averaged relaxation in the cluster.
The authors acknowledge some differences with respect to our work: (i) they report eccentricities at $r_{\mathrm{p}}=10\,R_{\mathrm{g}}$ rather than at the separatrix; (ii) they set $t_{\mathrm{GW}}$ to the approximate coalescence time from 2020MNRAS.495.2321Z, rather than the timescale over which GWs instantaneously impact the orbital elements that we adopt in 2025A&A...694A.272M [^11] We remark that the relevant timescale for determining whether angular momentum relaxation dominates over GW emission is the time needed by GWs to impact the current orbital elements, which is non-trivially related to the time it takes for the binary to merge.; (iii) they derive orbital parameters from the Newtonian potential of the MBH plus the stellar potential, while we use the map from 2024PhRvD.109f4077H. In Appendix D of their work, all three differences are removed, yielding distributions qualitatively similar to ours (see their Fig. 32 and our Fig. 5), but missing a large fraction of high eccentricity events, consistent with their non-inclusion of initially wide EMRIs. Once those EMRIs are removed from our dataset, the distributions become much more similar in shape, although theirs peaks at slightly higher eccentricity (see Fig. 9). This is expected, as Peters’ equations are known to overestimate eccentricity near plunge, where their underlying assumptions break down [2021PhRvD.104j4023T, 2025PhRvD.112b4012F].
[…]
Figure 9: EMRI eccentricity distribution at plunge for $M_{\bullet}=10^{5}\,\mathrm{M_{\odot}}$. The purple dashed histogram shows the time-averaged distribution from 2025arXiv250900469S, averaged up to $t=10\,t_{\mathrm{peak}}$ and obtained under assumptions similar to ours (see text). The green histogram shows our results for the classical EMRI population only. The data are the same as in Fig. 5, with the normalization adjusted accordingly.
Since a significant fraction of our initial conditions have $r_{\mathrm{p}}<10\,R_{\mathrm{g}}$, the direct comparison of the eccentricity distribution at $r_{\mathrm{p}}=10\,R_{\mathrm{g}}$ is practically unfeasible.
@@
Finally, 2017PhRvD..95j3012B reported a flat eccentricity distribution at plunge in the range $0<e_{\mathrm{pl}}<0.2$, with a small tail of outliers with higher $e_{\mathrm{pl}}$. This distribution was derived as a suitable proxy to a set of numerical integrations of EMRIs within the classical capture radius $a_{\mathrm{c}}$ to plunge. It should be noted, however, that although the $e_{\mathrm{pl}}$ information is given in their paper, it is not used at any point in their calculations, which focus on the computation of cosmological EMRI rates rather than on the determination of detailed EMRI orbital properties. Nevertheless, a flat $e_{\mathrm{pl}}$ distribution in the range $[0,0.2]$ has become a somewhat standard assumption when simulating EMRI populations. Our work demonstrates how inaccurate this assumption is (see Fig. 5), and highlights the much larger eccentricity range covered by EMRIs formed via two-body relaxation.
@@ V Conclusions
In this work we derived realistic EMRI eccentricity distributions around Schwarzschild MBHs assuming the two-body relaxation channel. We quote the distribution at different stages of the inspiral, up to the final plunge. We accomplished this by integrating further the simulations performed by 2025A&A...694A.272M using the FEW package, appropriately weighting each run proportionally to the EMRI formation rate at its initial semi-major axis.
@@
There are some effects that we neglected in this work, which could influence the eccentricity distribution of EMRIs. First, we only considered Schwarzschild BHs, whereas astrophysical BHs are generally expected to be spinning. The spin of the stellar-mass BH is unlikely to induce a significant deviation in the eccentricity of EMRIs at the level of precision considered here (the effect is small and enters at post-adiabatic order; see e.g. Ref. [2024PhRvD.109l4048B]). This, however, is likely not the case for the spin of the MBH. The loss of orbital energy and angular momentum differs for inspirals around Kerr MBHs, and, most importantly, the location of the separatrix depends on both the spin of the MBH and the inclination of the orbital plane with respect to the MBH spin axis [2020PhRvD.101f4007S]. From an astrophysical perspective, this affects EMRI formation rates: plunging directly into a highly spinning MBH is harder on prograde orbits and easier on retrograde ones, compared to the Schwarzschild case (spin effectively shrinks or enlarges the loss cone region). When considering a population of EMRIs with generic inclinations around a Kerr MBH, the overall effect is expected to be an enhancement in the EMRI formation rate [2013MNRAS.429.3155A]. We expect this effect to be more pronounced for initially wide orbits, thus increasing the number of highly eccentric EMRIs compared to the results presented here. We plan to properly investigate the impact of MBH spin in future work.
Another simplification adopted in this study is the assumption that all stellar-mass BHs have the same mass of $10\,\mathrm{M_{\odot}}$. A more realistic eccentricity distribution should account for a mass spectrum of the secondary objects and the corresponding mass segregation effects as the cluster evolves. We believe, however, that the impact of this would be small, since a continuous mass spectrum introduces a scatter in the slope of the distribution of the orbiting bodies comparable to the scatter observed during profile relaxation (compare our Fig. 3 to Fig. 10 in Ref. [2024ApJ...961..232Z]; our nuclear star cluster model has slope indices for stars and stellar-mass BHs close to their model M7). As discussed in Sect. II.2, this scatter does not strongly affect the eccentricity distribution. Even if segregation might have a minor impact, heavier objects might have different eccentricity distributions because of stronger GW emission. This should be investigated further, especially as evidence for a subpopulation of stellar-mass BHs around $35\,\mathrm{M_{\odot}}$ is emerging from GW observations [2025arXiv250818083T]. Nonetheless, 2005ApJ...629..362H showed that the eccentricity distribution is not strongly affected if the secondary is assumed to be a lighter compact object, such as a neutron star or a white dwarf, and we might expect a similar behavior for more massive objects.
Our simplified nuclear star cluster model also neglects the presence of binaries. Strong scatterings with tight binaries can lead to additional energy exchange due to the internal binding energy of the binary, influencing the relaxation process. As the average separation between objects in the nuclear star clusters we consider is much larger than typical binary separations ($0.1\,\text{-}\,100\,\mathrm{AU}$), the exchange of energy with the inner binary might be subleading. Nevertheless, this possibility should be explored in future work to confirm that the eccentricity distributions inferred here are robust.
Despite these caveats, our results call for an extension of the capabilities of FEW in order not to miss a large part of the EMRI population. We stress that any waveform model that aims at modeling astrophysical EMRIs has to cover a much wider parameter space that currently available. Accurate modeling of EMRIs is fundamental for parameter estimation, tests of general relativity and environmental effects, and for avoiding spurious residual in LISA global fit procedures. Precisely detecting high eccentricity EMRIs is also essential if we wish to infer the underling formation rate of each channel, which in turn will reveal more about the nature of MBHs and their host galaxies.
@@ Acknowledgements.
We thank Niels Warburton for useful discussions during the “Enabling Future Gravitational Wave Astrophysics in the Milli-Hertz Regime” workshop, as well as the organizers and administrative staff for ensuring a stimulating and well-run meeting. We also thank Monica Colpi for her valuable comments and suggestions.
D.M. acknowledges that this publication was produced while attending the PhD program in Space Science and Technology at the University of Trento, Cycle XXXIX, with the support of a scholarship financed by the Ministerial Decree no. 118 of 2nd March 2023, based on the NRRP - funded by the European Union - NextGenerationEU - Mission 4 “Education and Research”, Component 1 “Enhancement of the offer of educational services: from nurseries to universities” - Investment 4.1 “Extension of the number of research doctorates and innovative doctorates for public administration and cultural heritage” - CUP E66E23000110001. A.S. acknowledges the financial support provided under the European Union’s H2020 ERC Advanced Grant “PINGU” (Grant Agreement: 101142079). M.B. acknowledges support from the Italian Ministry for Universities and Research (MUR) program “Dipartimenti di Eccellenza 2023-2027”, within the framework of the activities of the Centro Bicocca di Cosmologia Quantitativa (BiCoQ). This research was supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP) which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – EXC-2094 – 390783311.
Data availability
The data that support the findings of Fig. 5 are openly available on Zenodo [dataZenodo].
@@ Appendix A Alternative weighting
In this appendix, we examine how the results in Fig. 5 change when using an alternative weighting procedure based on the classical criterion for EMRI formation. In the standard approach [2005ApJ...629..362H], BHs entering the loss cone are classified as EMRIs if $a<a_{\mathrm{c}}$, and as direct plunges if $a>a_{\mathrm{c}}$. By contrast, in Sect. II.2 we employed a transfer function $S$ to provide a smoother transition between the two regimes and to account for cliffhanger EMRIs.
@@
in the energy range explored in 2025A&A...694A.272M. We then re-process the same simulations as in Sect. II.2 to compute the formation rate of EMRIs according to the classical differential rate
@@
Figure 10: EMRI eccentricity distribution at plunge, for different central MBH masses. Here we report the probability density function of our EMRI population analyzed with the procedure described in Sect. II.2 (blue solid histograms) and the one described in this appendix (green dotted histograms). We also show (orange dashed histogram) the distribution reported in 2017PhRvD..95j3012B, which has become a standard assumption in the literature.